# Surrogate Neural Networks Local Stability for Aircraft Predictive Maintenance

### v2

@@ 1 Introduction
There have been significant advances in the development of verification methods assessing the robustness of neural networks (NN), in particular against perturbations, often represented by $l_{p}$ norm balls. (e.g, Bak et al. (2020); Ehlers (2017); Balunovic et al. (2019); Botoeva et al. (2020b); Bunel et al. (2020); Henriksen and Lomuscio (2020, 2021); Katz et al. (2017, 2019); Kouvaros and Lomuscio (2021)).
Formal methods are now seen as key approaches to meet the emerging certification requirements for NN use in critical systems e.g., in aviation. Notable references, including guidelines from certification and aviation standardization entities, are advocating for the use of formal methods as a mean of compliance during the learning process management and inference model verification phases (see e.g., the European Aviation Safety Agency (EASA) recent concept paper ‘First usable guidance for Level 1 machine learning applications [^1] https://www.easa.europa.eu/en/easa-concept-paper-first-usable-guidance-level-1-machine-learning-applications-proposed-issue-01pdf ). A concrete example of this use is introduced in the ForMuLA report (Formal Methods use for Learning Assurance), published from the partnership between EASA and Collins Aerospace (ForMuLA (2023)). It includes an application of formal methods to an industrial use case of prediction of the remaining life of aeronautical components and illustrates how these verification techniques can be used to assess the safety of machine-learning (ML) based systems.
@@
In this paper, we are conducting the robustness assessment of NN surrogate models designed in the context of civil aviation and predictive aircraft maintenance. The case study presents unique specificities (e.g., regression task, multiple inputs and outputs etc.), rarely explored in the current literature. In this case study, we explore the adaptability and scalability of several families of verification techniques to the use case at hand. We further build upon their strengths and limitations to optimise their sequential and complementary use in order to perform the complete NN robustness assessment on a representative test set.
Section 2 introduces the case study of aircraft loads-to-stress prediction along with the tested models and stability property to be ensured. Section 3 describes the combination via the sequential use of techniques we employ in order to perform the NN stability assessment. Section 4 details the experiments and illustrates the verification process on a few NN models. Sections 5 and 6 present findings and conclusions.
@@ 2.1 Description
In aeronautics, numerical simulations are regularly used to model complex physical phenomena in systems or structures (using, e.g., a finite element discretization approach) and understand and anticipate their behaviour. The computational cost of such simulations has however prevented their use in real-time, including their embeddability in products or on-line processes. Their application therefore remains limited e.g., to a system design phase. The use of such simulation could however greatly improve productivity and operational cost, e.g., in the domain of aircraft predictive maintenance.
Structural maintenance programs (e.g., replacement of parts etc.) are currently based on conservative assumptions: they often assume aircrafts of a fleet have sustained a similar flight history and, thus, operational fatigue, and adopt worse case scenario, often overly preventive, maintenance actions for the whole fleet. One can therefore see the benefit in introducing a more optimised maintenance program, tailored to individual aircraft in order to reduce individual maintenance cost while ensuring that the safety of operations is maintained at the highest level. Such custom solutions, however, require the analysis of large amount of operational data from flight past history, the complete modelisation of the sustained aircraft fatigue and the prediction of the impact it had on its different parts in order to anticipate the required maintenance operations. These simulations involve computationally-demanding physics and mathematical modeling that may be unfeasible to use in real-life operational settings.
NN are a game changer. More and more studies are showing the value of NN-based surrogates to approximate numerical simulators (e.g., Sudakov et al. (2019)). Their more systematic use in industry, especially as components or enablers of safety critical applications however requires the maturation of new verification processes. In this analysis, we investigate an example of NN surrogate models trained to predict the level of stress in different parts of an aircraft structure from sustained external loads. More details on the model function is provided in O’Higgins et al. (2020). The accurate evaluation of the stresses is a key enabler for maintenance optimization.
@@ 2.2 Tested models
The NN tested in the present study are research prototypes and less mature versions of the models currently pushed at Airbus. The models that will be used in predictive maintenance settings are currently being optimised, in part supported by robustness analysis such as the one presented here.
@@ 2.3 Property to be ensured: Local stability
  • $\underline{\delta_{i}}$ and $\overline{\delta_{i}}$, upper and lower constrains on deviation from original prediction
  • $T$, a threshold defining two zones of the output space with distinct requirements on stability i.e., $|f_{i}(x)|\leq T$ (‘Knot’) and $|f_{i}(x)|>T$ (‘Wings’)
  • $c_{T}$, the set constraint on $\delta_{i}$ in the knot
  • $p_{out}$, the constrains on output perturbation in the wings
@@
This stability requirement is initially set to be very permissive; this will be illustrated later in the paper when the tested NN will show non-stability for a significant fraction of the test points (see §5). The tested models and property offer a nice playground for the benchmarking of verification methods.
@@ 3 Method combination for local stability assessment
Before the advent of formal NN verification methods, the model local stability was assessed at Airbus via a random sampling of its immediate ‘neighborhood’ and corresponding predictions. This ‘brute force’ approach was both partial and computational expensive due to the necessity to homogeneously cover via sampling a high number of perturbed input variables. There was therefore a crucial need for formal, sound and time-efficient means to assess the NN stability.
@@
Exploring a combination of techniques and using their strengths and limitations to optimise the verification assessment seems like a promising direction to fulfill the requirements of soundness and time-efficiency. We converge towards the sequential use of the following techniques:
@@
We conduct experiments generating several types of classical adversarial attacks e.g., Projected Gradient Descent approach (PGD; Madry et al. (2017)) or Fast Gradient Sign Method (FGSM; Goodfellow et al. (2014)). We make use of the cleverhans library [^2] https://github.com/cleverhans-lab/cleverhans. The goal of this first step is to quickly find examples of model local instability. We will later see that a simple attack generation technique such as PGD is sufficient to determine all test points whose local stability property is violated. We do not discard that more advanced attack generation techniques might be required as the models become more locally stable (e.g., via local stability training).
@@
We then use formal methods. We first leverage the less computationally-demanding nature of incomplete methods. We make use of the linear relaxation-based perturbation analysis (LiRPA) verification bound method CROWN (Zhang et al. (2018)) implemented within the Airbus open-source decomon [^3] https://github.com/airbus/decomon library (Ducoffe (2021)). CROWN is a state-of-the-art technique in the field of abstract interpretation and has been shown to provide tighter bounds compared to other techniques, which enhances its effectiveness for neural network verification.
@@
To evaluate test points whose stability is neither refuted nor guaranteed by the previous methods, we use an in-house Mixed-Integer Linear Program (MILP)-based verifier. Encoding of the network into variables and constraints is done similarly to the Venus library (Botoeva et al. (2020a)). Symbolic interval propagation (Wang et al. (2018)) is used to provide a first bound evaluation for neuron values. Details on the implementation of the MILP-based solution are provided in Appendix A.
The sequence of these verification techniques is illustrated in Fig. 1. While we evaluate the efficiency of this verification process via this specific set of techniques (see §4), we expect that similar analysis can be conducted by removing bricks, sequentially combining techniques in another order (e.g., B+A+C instead of A+B+C) or replacing each technological brick by any technique falling into the same family (incomplete by incomplete etc.).
@@ 4 Experiments
The assessment of local stability (§2) is performed on $1000$ test points whose distribution matches the one of the training set and is representative of the operating domain.
@@ A B C A+C B+C Pipeline A+B+C
Table 1 summarises the stability assessment performed on two of the models ($2$ and $500$-epoch trained for $2$ and $500$ epochs respectively) using the verification pipeline. Attacks are generated using the PGD technique using $20$ iterations of step size $\epsilon=0.01$. The MILP verifier is run with a timeout $t=300$s per point. None of the test points’ verification exceeded this timeout.
As stated earlier, adversarial attack generation can only provide ‘False’ status, but cannot guarantee the property is ‘True’. Incomplete/CROWN, on the opposite, can only prove that the property is ‘True’ for some test points i.e., they guarantee stability but not its lack of. Only complete techniques can unequivocally prove that a point stability is either guaranteed or disproved.
@@ 5.1 A significant verification time gain
Similarly, the three-stage pipeline has a significantly faster runtime than a portion of it (e.g., Bricks A+C) which further shows the effectiveness of combining techniques, thereby reducing the call to time-demanding exact computations. The use of incomplete methods as Brick B allows e.g., to converge towards a guarantee of local stability for the large majority of the test data that were still to be evaluated after the adversarial attack stage (e.g., $554/558$ and $529/552$ for $2$-epoch and $500$-epoch models respectively).
A subtlety of the complete MILP-based method implemented in the present analysis is that it uses bound propagation (Wang et al. (2018)) in order to provide lower and upper values to neurons in the network. When the bound propagation already allow to reach a definite conclusion on the stability property status, there is no call to the MILP solver (see Fig. 2). An example can be seen in the time of method A+C for the 2-epoch model competitive with respect to the full pipeline (Table 1, column $4$ vs. column $6$). On this lightly trained model, the first step of bound propagation estimation already concludes on the stability of all but four test points, without the need for additional MILP computations. On the contrary, when using method B+C (column 5), most of the points entering the complete method are not locally stable. They would have been identified during the adversarial attack phase (not used in B+C) and the MILP-based solver has to be called to find or not counter-examples to the property.
@@ 5.2 Insights into the models
We first observe that some output indexes are more vulnerable than others, a result that is carefully being considered for the on-going model robustification. In the present analysis, the adversarial attacks are performed in order of output indexes. We however learned, e.g., from the adversarial attack experiments, that some indexes are more easily attackable than others. We can benefit greatly from this knowledge to optimise the attack generation on subsequent generation of models and gain time on future robustness assessment.
@@
6 Conclusions and ways forward
This case study analysis presents the implementation and testing of a sound multi-technique verification pipeline for the a posteriori assessment of the local stability property of surrogate NN. By combining the strengths and mitigating the limitations of a number of techniques including adversarial attacks and complete and incomplete formal verification methods, we have not only managed to successfully perform a complete NN stability assessment of the models over the test dataset but also reduced the runtime of the required experiments to do so. Furthermore, the experiments have also shone light on model vulnerabilities as well as the necessity to investigate the feasibility of looser constraints for the local stability property itself.
This case study and technique configuration only begins to explore what such a verification pipeline can offer; further tuning of the techniques could indeed lead to an even more time-efficient verification. We performed a number of optimisation experiments e.g., hyperparameters tuning for adversarial attacks, MILP-computation etc. But further optimisation work could be done in that respect. As far as the implementation of the MILP solver is concerned (see Appendix A), the present analysis uses symbolic interval propagation in order to provide starting bounds and restrict the input domain to be verified. For more complex NN however, interval arithmetic could provide excessively conservative bounds that may slow the verification. We will be exploring the use of bounds derived from an incomplete verification tool (e.g., CROWN) in order to feed them as initial bounds to the complete verifier.
The next step is the enhancement of the NN surrogate models through the integration of empirical and formal components as part of the model design phase (e.g., during training), moving beyond the sole a posteriori evaluation of its local stability.

### v3

@@ 1 Introduction
There have been significant advances in the development of verification methods to assess the robustness of neural networks (NN). Formal methods are, in particular, considered as key approaches to be explored and matured (see §2). They are expected to provide model guarantees that could allow industries, such as aviation, to meet the emerging certification requirements for the use of NN, in particular for safety-critical systems. Notable references, including guidelines from certification and aviation standardization entities, are advocating for the use of formal methods as a mean of compliance during the learning process management and inference model verification phases (see, e.g. the European Aviation Safety Agency (EASA) recent concept paper ‘First usable guidance for Level 1 machine learning applications [^1] https://www.easa.europa.eu/en/easa-concept-paper-first-usable-guidance-level-1-machine-learning-applications-proposed-issue-01pdf ). A concrete example of this use is introduced in the ForMuLA report (Formal Methods use for Learning Assurance), published from the partnership between EASA and Collins Aerospace ForMuLA (2023). It includes an application of formal methods to an industrial use case of prediction of the remaining life of aeronautical components and illustrates how these techniques can be used to assess the safety of machine-learning (ML) based systems.
@@
In this paper, we are conducting the robustness assessment of NN surrogate models designed in the context of civil aviation and predictive aircraft maintenance. The case study presents unique specificities (e.g. regression task, multiple inputs and outputs etc.), rarely explored in the current literature. In this case study, we explore the adaptability and scalability of several families of verification techniques to the use case at hand. We further build upon their strengths and limitations to optimise their sequential and complementary use in order to perform the complete NN robustness assessment on a representative test set.
Section 2 summarizes the recent advances in NN verification. Section 3 introduces the case study of aircraft loads-to-stress prediction along with the tested models and stability property to be ensured. Section 4 describes the combination via the sequential use of techniques we employ in order to perform the NN stability assessment. Section 5 details the experiments and illustrates the verification process on a few NN models. Sections 6 and 7 present findings and conclusions.
2 Related Work
NN are brittle. The emergence of the field of adversarial machine-learning has shown that models can be easily fooled, even by small perturbations to input data Goodfellow et al. (2014). A series of work has subsequently attempted to craft defense mechanisms against adversarial attacks e.g., madry2019deep. Ultimately, however, these techniques do not provide guarantee that a neural network is robust to perturbations but only that counterexamples to a robustness requirement to be guaranteed can be generated.
The introduction of adversarial attacks has motivated the ramp-up of the applicability of formal methods to AI models, in particular to assess the robustness of neural networks against small ‘local’ perturbations, often represented by $l_{p}$ norm balls. Two families of verification methods have since been matured: (i) incomplete techniques e.g., Katz et al. (2017, 2019); Kouvaros and Lomuscio (2021) that can prove the presence/absence of counterexamples to a given property. They are usually computationally expensive and scale poorly with large networks. (ii) complete techniques, e.g., NEURIPS2018_f2f44698; DBLP:journals/pacmpl/SinghGPV19; Zhang et al. (2018); xu2021fast, that provide approximated bounds to model prediction. They commonly perform faster and thus scale better with larger models. They can however severely over approximate bound estimates of model output, leading to their lack of convergence for a given robustness status.
Verification tools are now commonly combining complete and incomplete techniques wu2024marabou in order to speed the verification process. However, to the best of our knowledge, there has been very limited published works illustrating and quantifying the impact (e.g. on computing time) this combination brings. As model size and input dimensionality keep increasing, scalability and verification optimisation will however be of essence for the adoption of formal methods in the verification, validation and certification processes of AI-based industrial systems.
@@ 3.1 Description
In aeronautics, numerical simulations are regularly used to model complex physical phenomena in systems or structures (using, e.g. a finite element discretization approach) and understand and anticipate their behaviour. The computational cost of such simulations has however prevented their use in real-time, including their embeddability in products or on-line processes. Their application therefore remains limited e.g. to a system design phase. The use of such simulation could however greatly improve productivity and operational cost, e.g. in the domain of aircraft predictive maintenance.
Structural maintenance programs (e.g. replacement of parts etc.) are currently based on conservative assumptions: they often assume aircrafts of a fleet have sustained a similar flight history and, thus, operational fatigue, and adopt worse case scenario, often overly preventive, maintenance actions for the whole fleet. One can therefore see the benefit in introducing a more optimised maintenance program, tailored to individual aircraft in order to reduce individual maintenance cost while ensuring that the safety of operations is maintained at the highest level. Such custom solutions, however, require the analysis of large amount of operational data from flight past history, the complete modelisation of the sustained aircraft fatigue and the prediction of the impact it had on its different parts in order to anticipate the required maintenance operations. These simulations involve computationally-demanding physics and mathematical modeling that may be unfeasible to use in real-life operational settings.
NN are a game changer. More and more studies are showing the value of NN-based surrogates to approximate numerical simulators Sudakov et al. (2019). Their more systematic use in industry however requires the maturation of new verification processes, specific to machine-learning components. The current lack of consensus on the tools to conduct such verification dramatically limits their applications, especially as components or enablers of safety critical applications. In this analysis, we investigate an example of NN surrogate models trained to predict the level of stress in different parts of an aircraft structure from sustained external loads. More details on the model function is provided in O’Higgins et al. (2020). The accurate evaluation of the stresses is a key enabler for maintenance optimization.
@@ 3.2 Tested models
The NN tested in the present study are research prototypes and earlier, less mature versions of the models currently pushed at Airbus. The models that will be used in predictive maintenance settings are currently being optimised, in part supported by robustness analysis such as the one presented here.
@@ 3.3 Property to be ensured: Local stability
  • $\underline{\delta_{i}}$ and $\overline{\delta_{i}}$ are respectively upper and lower constrains on deviation from original prediction,
  • $T$ is a threshold defining two zones of the output space with distinct requirements on stability i.e., $|f_{i}(x)|\leq T$ (‘Knot’) and $|f_{i}(x)|>T$ (‘Wings’),
  • $c_{T}$ is the set constraint on $\delta_{i}$ in the knot and
  • $p_{out}$ are constrains on output perturbation in the wings.
@@ 4 Method combination for local stability assessment
Before the advent of formal NN verification methods, the model local stability was assessed at Airbus via a random sampling of its immediate ‘neighborhood’ and corresponding predictions. This ‘brute force’ approach is both partial and computational expensive due to the necessity to homogeneously cover via sampling a high number of perturbed input variables. There was therefore a crucial need for formal, sound and time-efficient means to assess the NN stability.
@@
Exploring a combination of techniques and using their strengths and limitations to optimise the verification assessment seemed like a promising direction to fulfill the requirements of soundness and time-efficiency. We converge towards the sequential use of the following techniques:
@@
We conduct experiments generating several types of classical adversarial attacks, e.g. Projected Gradient Descent approach (PGD) Madry et al. (2017) or Fast Gradient Sign Method (FGSM) Goodfellow et al. (2014) specifically designed to find local violation of model stability. We make use of the cleverhans library [^2] https://github.com/cleverhans-lab/cleverhans and adapt the adversarial loss function to the given local stability property. The goal of this step A is to quickly find examples of model local instability. We notice that a simple attack generation technique such as PGD is sufficient to determine all test points whose local stability property is violated. We do not discard that more advanced attack generation techniques might be required as the models become more locally stable (e.g. via local stability training). The stability of test points for which no attack was generated is assessed using formal methods.
@@
We leverage the less computationally-demanding nature of incomplete methods in order to assess test points whose stability might be less challenging to guarantee. We intend to provide lower and upper bound estimates of the model outputs to perturbed inputs. To do so, we make use of the linear relaxation-based perturbation analysis (LiRPA) verification bound method CROWN Zhang et al. (2018).
CROWN (or its equivalent DeepPoly DBLP:journals/pacmpl/SinghGPV19) is a commonly adopted bound propagation method. It has been shown to provide tighter bounds compared to previously developed techniques by means of linear and quadratic functions which enhance its effectiveness on many activation functions including ReLUs (used in our models). We make use of the version of CROWN implemented within the Airbus open-source decomon [^3] https://github.com/airbus/decomon library Ducoffe (2021).
@@
To evaluate test points whose stability is neither refuted nor guaranteed by the previous methods, we use an in-house Mixed-Integer Linear Program (MILP)-based verifier. Encoding of the network into variables and constraints is done similarly to the Venus library Botoeva et al. (2020a).
The MILP encoding makes use of bounds computed by symbolic interval propagation Wang et al. (2018). For each layer $i$ and neuron $j$ in the layer $i$, two continuous variables are created : $y_{i,j,-},y_{i,j,+}$ which corresponds to pre-activation and post-activation values of the neurons. Bounds of these variables are known thanks to the bound propagation function. Layer $i=0$ corresponds to input values. The MILP is encoded as follow:
  1. 1.
  2. 2.
The $max$ constraint is native into the Gurobi library and can also be chosen instead of the classically used Big-M constraint Botoeva et al. (2020a).
The negation of the stability property is encoded as follows. Given $x,f(x),\underline{\delta},\overline{\delta}$ defined in property 1:
  1. 1.
    Let $\forall j\in[1,81],indic_{-,j},indic_{+,j}$ be binary variables.
  2. 2.
  3. 3.
  4. 4.
    $\sum_{j}{(indic_{+,j}+indic_{-,j})}\geq 1$
Each indicator variable will encode the fact that, for one output index, the perturbed NN prediction goes out of the bow-tie property. The last constraint thus encodes that the solver has to find a counterexample violating the property. If the MILP solver finds a solution, we conclude that the property is ‘False’. If the solver shows the absence of counterexample, we conclude that the property is ‘True’.
The sequence of these verification techniques is illustrated in Fig. 1. While we evaluate the efficiency of this verification process via this specific set of techniques (see §5), we expect that similar analysis can be conducted by removing bricks, sequentially combining techniques in another order (e.g. B+A+C instead of A+B+C) or replacing each brick by any technique falling into the same family (incomplete by incomplete etc.).
In order for the community to be able to perform similar verification technique combination and benchmarking, Airbus has made the verification pipeline open-source. Link and details on the source code are provided in Appendix A.
@@ 5 Experiments
The assessment of local stability (§3) is performed on $1000$ test points [^4] We acknowledge that the present techniques allow the verification of the local stability property of a finite number of points in the input domain and do not permit in their current implementation to prove the global robustness of a continuous input domain. Work is currently being pushed by the formal method community towards exploring the extension of these techniques to global robustness verification. whose distribution matches the one of the training set and is representative of the operating domain.
@@
Attacks are generated using the PGD technique using $20$ iterations of step size $\epsilon=0.01$. The MILP verifier is run with a timeout $t=300$s per point. None of the test points’ verification exceeded this timeout.
Table 1 summarises the stability assessment performed on two of the models (trained with $2$ and $500$ epochs) using the verification pipeline. As stated earlier, adversarial attack generation can only provide ‘False’ status, but cannot guarantee the property is ‘True’. Incomplete/CROWN, on the opposite, can only prove that the property is ‘True’ for some test points i.e., they guarantee stability but not its lack of. Only complete techniques can unequivocally prove that a point stability is either guaranteed or disproved.
@@ A B C A+C B+C Pipeline A+B+C
We see that we can generate attacks for $442$ ($448$) out of $1000$ test points for the $2$-epoch ($500$-epoch) model. We can provide local stability guarantees for $554$ ($529$) of the remaining $558$ ($552$) test points. When combining bricks A+B+C, we only have to perform complete verification on the last $4$ ($23$) test points.
@@ 6.1 A significant verification time gain
Similarly, the three-stage pipeline has a significantly faster runtime than a portion of it (e.g. Bricks A+C) which further shows the effectiveness of combining techniques, thereby reducing the call to time-demanding exact computations. The use of incomplete methods as Brick B allows, e.g. to converge towards a guarantee of local stability for the large majority of the test data that were still to be evaluated after the adversarial attack stage (e.g. $554/558$ and $529/552$ for $2$-epoch and $500$-epoch models respectively).
A subtlety of the complete MILP-based method implemented in the present analysis is that it uses bound propagation Wang et al. (2018) in order to provide lower and upper values to neurons in the network. When the bound propagation already allow to reach a definite conclusion on the stability property status, there is no call to the MILP solver (see Fig. 2). We acknowledge that that the derivation of these bounds are conceptually similar to bounds derived using incomplete methods. While this part of brick C is here performed using symbolic interval analysis that provides tight but potentially looser bounds than CROWD, we are exploring the feasibility of using bounds derived by method B as inputs to C.
An example can be seen in the time of method A+C for the 2-epoch model competitive with respect to the full pipeline (Table 1, column $4$ vs. column $6$). On this lightly trained model, the first step of bound propagation estimation already concludes on the stability of all but four test points, without the need for additional MILP computations. On the contrary, when using method B+C (column 5), most of the points entering the complete method are not locally stable. They would have been identified during the adversarial attack phase (not used in B+C) and the MILP-based solver has to be called to find or not counter-examples to the property.
@@ 6.2 Insights into the models
We first observe that some output indexes are more vulnerable than others, a result that is being used for the on-going model strengthening. Indeed, if the local stability is disproved for at least one output index, the test point is considered as not meeting the safety requirement. If some indexes are systematically shown to be vulnerable, the pertinence of these indexes might be re-evaluated.
In this analysis, the adversarial attacks are performed in order of output indexes. We can also greatly benefit from the fact that we can anticipate which indexes are more easily attackable than others to optimise the attack generation on subsequent generation of models and gain time on future robustness assessment.
@@
7 Conclusions and future work
This case study analysis presents the implementation of a sound multi-technique pipeline for the a posteriori assessment of the local stability property of surrogate NN. By combining the strengths and mitigating the limitations of a number of techniques including adversarial attacks and complete and incomplete formal verification methods, we have not only managed to successfully perform a complete NN stability assessment of the models over the test dataset but also reduced the runtime of the required experiments to do so. Furthermore, the experiments have also shone light on model vulnerabilities and the necessity to investigate the feasibility of looser constraints for the local stability property itself.
This case study and technique configuration only begins to explore what such a verification pipeline can offer; further tuning of the techniques could indeed lead to an even more time-efficient verification. We performed a number of optimisation experiments e.g., hyperparameters tuning for adversarial attacks, MILP-computation etc. But further optimisation work could be done in that respect. As far as the implementation of the MILP solver is concerned (see §4), the present analysis uses symbolic interval propagation in order to provide starting bounds and restrict the input domain to be verified. For more complex NN however, interval arithmetic could provide excessively conservative bounds that may slow the verification. We will be exploring the use of bounds derived from an incomplete verification tool (e.g. CROWN) in order to feed them as initial bounds to the complete verifier.
The next step is to enhance NN surrogate models through the integration of empirical and formal components as part of the model design phase (e.g. during training), moving beyond the sole a posteriori evaluation of its local stability.